### v3

@@ 4.3 Explainability Component
Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data. It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach when understanding the behavior of autonomous robots.
4.4 Software and Hardware Artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
The chosen ROS 2 version is Humble. Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table 2. These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging purposes.
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [87]. This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
The interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), and ChromaDB (version 0.4.6). We used the open-source model Llama 2 7B GGUF [88] as the LLM responsible for generating natural language explanations for end users. This LLM integrates Reinforcement Learning from Human Feedback (RLHF) and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral [89] or Zephyr [90], led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase. As a text embedding model we have chosen Instructor-xl [91], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, a fact particularly important in the processes performed for any RAG solution.
Table 2: Rosbag topics rate and average rate per message.
@@ 4.4 Software and Hardware Artifacts
5 Evaluation
To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4. This study covered several approaches for storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as a source the data gathered by our black box-like component, through a comprehensive set of questions. This work aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
@@ 5 Evaluation
This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people. Details like the initial position and the number of waypoints the robot needs to visit are provided beforehand. Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during runtime. Each waypoint is characterized by its $X$, $Y$, and $\Theta$ coordinates. Existing elements in the environment, such as furniture, doors, or walls, must be mapped. The obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its previously planned path to reach the second goal. This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal. Consequently, this obstacle prompts the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as detailed in the second scenario.
The three scenarios are represented in Figure 4.
The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and is licensed under GPLv3.
@@
5.1 Black Box Recorder Evaluation
In evaluating the accountability component, we conducted six experiments across three previously described scenarios. These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@ 5.1.1 Metrics
We employed different metrics to evaluate the impact of the accountability system in the chosen scenarios. First, through the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the employed blockchain storage approach.
@@
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure. This information is included in Section 6.
5.2 Explainability Evaluation
@@
Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have led to emerging trends in LLM evaluations, such as the use of strong LLMs as judges [82].
@@
Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. Furthermore, this approach significantly impacts the scalability of evaluations, making it necessary to use more effective and unbiased methods [81, 93]. To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4. These models are often trained with RLHF, resulting in strong human alignment [94]. Previous studies have evidenced the consistent agreement of these models with human grading [95, 96, 97, 98]. Moreover, research shows that LLMs can replicate most of conclusions from human evaluations, particularly in tasks involving prediction accuracy and explanation helpfulness, even across varied settings and domains [81, 99, 100].
@@
However, while LLMs excel in scalability and efficiency, their strengths can be enhanced by integrating the nuanced understanding and subjective judgments that humans provide, particularly in complex or domain-specific contexts [101]. For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method might overlook [48, 102].
@@
To address these limitations, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessments. In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessments. Human evaluators then validate a representative subset of these outputs to uncover overlooked biases or misalignments. This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly for nuanced cases.
@@
5.2.1 LLM-as-a-judge Evaluation
The need for a robust LLM aligned with human preferences led us to select GPT-4 to evaluate the answers produced by our explainability component. Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
@@ 5.2.1 LLM-as-a-judge Evaluation
LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times. This behavior derives from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response. The effect of this indeterminism must be considered in any LLM evaluation.
@@
In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims to determine the relevance of its document selection concerning a query. To this end, we conducted an experimental study with 16 different questions on the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
@@
Using the context provided by the Rosbag data and an evaluation question, our approach includes an evaluation component for the generated explanations that contrasts the prediction made by the evaluated LLM and, thanks to the retrieved selection, compares it to an annotated ground truth. This process yields a correctness metric that scores whether the retriever returned no relevant documents (0) or whether it returned a perfect selection of documents (1). The final score, an average of all the scores in the dataset, is used to judge the retriever’s overall performance.
@@
A comprehensive evaluation of an LLM should consider multiple metrics that include correctness as well as other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. Produced responses must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable way to the user.
@@
Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are described in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
Table 3: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
@@
Accuracy Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
Helpfulness Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
Conciseness Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
Coherence Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
{mdframed} [hidealllines=true,backgroundcolor=blue!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt, skipabove=-3pt,skipbelow=-10pt] Question: How many goals have been reached by the robot? {mdframed}[hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals. The context mentions that the goal number 1, 2 and 3 have been reached successfully. Therefore, the answer to your question is 3.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful. It directly answers the user’s question by confirming that the robot has reached three goals. The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness. There are no unnecessary words or unclear statements, making the response concise and coherent. Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth. The response is concise, with no unnecessary words or information. It is highly coherent and has a clear structure. The response is extremely helpful as it gives the exact information asked for by the user. Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals. The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded. Position: -9.14068582195674, -25.977117025927537. Orientation: -0.6580310307304338,0.7529908117605705. * Goal number 3: Navigation to the goal number 3 has succeeded. Position: -1.8811226231679194, -28.773672979732826. Orientation: -0.5946646510401259,0.803973850820613. The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted. Position: -5.062160931312302, -8.330878461767831. Orientation: -0.8760794639388502,0.4821667479872674. Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals. Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate. It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted. The assistant also provides additional details about the positions and orientations of the goals. While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding. Therefore, the assistant’s response is also helpful. However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation. The assistant’s response is coherent as it logically organizes the information about each goal. Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0

### v4

@@ IV-C Explainability Component
Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data. It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach for understanding the behavior of autonomous robots.
IV-D Software and Hardware Artifacts
Our solution has been deployed on a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
The chosen ROS 2 version is Humble. Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table II. These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging.
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotic simulations. To provide a realistic scenario, we used a hospital simulation environment made available by the Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [81]. This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
Interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), and ChromaDB (version 0.4.6). We used the open-source model Llama 2 7B GGUF [82] as the LLM responsible for generating natural language explanations for end users. This LLM integrates Reinforcement Learning from Human Feedback (RLHF) and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral [83] or Zephyr [84], led us to choose this model for evaluating our data due to its promising results and suitability for the available hardware during the development phase of this work. As a text embedding model we have chosen Instructor-xl [85], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, which is particularly important in the processes performed for any RAG solution.
TABLE II: Rosbag topics rate and average rate per message.
@@ IV-D Software and Hardware Artifacts
V Evaluation
To evaluate our proposal, we conducted an experimental study to compare the performance of an accountability system built in accordance with Section IV. This study covered several approaches to storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using the data gathered by our black box-like component as a source, through a comprehensive set of questions. This study aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
@@ V Evaluation
This functionality evaluates the capability of a robot to navigate accurately and autonomously within a given scenario, including furniture, objects, and people. Details, such as the initial position and the number of waypoints the robot needs to visit, are provided beforehand. Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during the runtime. Each waypoint is characterized by its $X$, $Y$, and $\Theta$ coordinates. Existing elements in the environment, such as furniture, doors, and walls, must be mapped. Obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its previously planned path to reach the second goal. This adjustment becomes necessary because the appearance of an obstacle obstructing the original planned route, which requires a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door required to reach the first goal. Consequently, this obstacle leads to the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as those detailed in the second scenario.
The three scenarios are illustrated in Figure 4.
The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and licensed under the GPLv3.
@@
V-A Black Box Recorder Evaluation
To evaluate the accountability component, we conducted six experiments using the three previously described scenarios. These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one for every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@ V-A1 Metrics
We employed different metrics to evaluate the impact of the accountability system in the chosen scenarios. First, using the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the blockchain storage approach.
@@
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence a robot’s behavior and event-recording capabilities. In particular, we obtained a full report on system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure. This information is presented in Section VI.
V-B Explainability Evaluation
@@
Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, particularly in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have led to emerging trends in LLM evaluations, such as the use of strong LLMs as judges [22].
@@
Traditional similarity-based metrics, such as BLEU and ROUGE, have been proven to be misaligned with human evaluations. Although human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. Furthermore, this approach significantly affects the scalability of evaluations, making it necessary to use more effective and unbiased methods [21, 87]. To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs, such as GPT-4. These models are often trained with RLHF, resulting in strong human alignment [88]. Previous studies have shown the consistent agreement between these models and human grading [89, 90, 91]. Moreover, research shows that LLMs can replicate most of conclusions from human evaluations, particularly in tasks involving prediction accuracy and explanation helpfulness, even across varied settings and domains [21, 92, 93].
@@
However, while LLMs excel in scalability and efficiency, their strengths can be enhanced by integrating the nuanced understanding and subjective judgments that humans provide, particularly in complex or domain-specific contexts [94]. For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method may overlook [20, 95].
@@
To leverage the strengths of both approaches, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessment. In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessment. Human evaluators then validated a representative subset of these outputs to uncover overlooked biases or misalignments. This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly in nuanced cases.
@@
V-B1 LLM-as-a-judge Evaluation
The need for a robust LLM aligned with human preferences led us to select the GPT-4 model to evaluate the answers produced by our explainability component. Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable. The entire process is illustrated in Figure 5.
@@ V-B1 LLM-as-a-judge Evaluation
LLMs exhibit a certain degree of indeterminism, defined as a model’s ability to offer different answers when asked the same question or prompt several times. This behavior is derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, thereby introducing randomness to their responses. The effect of indeterminism must be considered in LLM evaluations.
@@
In addition, the use of RAG for question-answering systems is one of the most difficult LLM use cases for evaluation. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims to determine the relevance of document selection for a query. To this end, we conducted an experimental study with 16 different questions on the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times for each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, demonstrating the potential strengths and weaknesses of our solution.
@@
Using the context provided by the Rosbag data and a given evaluation question, our approach includes an evaluation component that assesses the generated explanations by comparing the prediction made by the LLM against an annotated ground truth. The comparison yields a correctness score of 1 if the prediction matches the ground truth, and 0 otherwise. This binary metric reflects the extent to which retrieved information supports accurate explanations. The final score, computed as the average across the dataset, measures the retriever’s overall effectiveness in providing context that leads to correct, grounded responses.
@@
A comprehensive evaluation of an LLM should consider multiple metrics including correctness and other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. The responses must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable manner to the user.
@@
Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation of accuracy, conciseness, coherence, and helpfulness. The selected criteria metrics are described in Table III. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the scenarios used. By employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of LLM output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three scenarios is shown in Figure 6.
TABLE III: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
@@
Accuracy Measures how well the response aligns with the reference, ranging from completely inaccurate to fully accurate.
Helpfulness Assesses the informational value of the response, from uninformative to highly insightful and comprehensive.
Conciseness Evaluates the brevity and relevance of the response, from overly verbose to succinct and focused.
Coherence Judges the logical structure and flow, from disorganized or inconsistent to well-structured and logically connected.